# Friederike Becker1, Fabian Krüger1 and Melanie Schienle1,2

### v3

@@ B.5.2 AR-annual and ARX-annual
$y_{t}=\nu+\alpha\leavevmode\nobreak\ y_{t-1}+u_{t},$
where $y_{t}$ is the annual observation (GDP growth rate or inflation rate) of interest, $\nu$ is an intercept term, $-1<\alpha<1$ is an autoregressive term and $u_{t}$ is the model’s error term that is assumed to be independently normally distributed with variance $\sigma^{2}$. We estimate $\nu$ and $\alpha$ via ordinary least squares and $\sigma^{2}$ by the mean of squared residuals.
To describe ARX-annual, let $\hat{y}_{\text{IMF},t,\text{FC}}$ denote the IMF point forecast for target year $t$ and target variable $y$, issued in fall of the same year (forecast horizon ”Fall, Current”). Similarly, $\hat{y}_{\text{IMF},t,\text{SC}}$, $\hat{y}_{\text{IMF},t,\text{FN}}$ and $\hat{y}_{\text{IMF},t,\text{SN}}$ respectively denote the point forecast for target year $t$, made in spring of the same year (forecast horizon ”Spring, Current”), fall of the previous year (forecast horizon ”Fall, Next”) and spring of the previous year (forecast horizon ”Spring, Next”). Note thus again that $t$ refers to the target year: for instance, $\hat{y}_{\text{IMF},t,\text{FC}}$ is issued during the same year (year $t$), while $\hat{y}_{\text{IMF},t,\text{FN}}$ is issued in year $t-1$. Both issue forecasts for the same target year.
@@
$y_{t}=\nu+\alpha y_{t-1}+\gamma\leavevmode\nobreak\ \hat{y}_{\text{IMF},t,\text{FC}}+u_{t},$
@@
$y_{t}=\nu+\alpha y_{t-1}+\gamma\leavevmode\nobreak\ \hat{y}_{\text{IMF},t,\text{SC}}+u_{t}.$
For next-year forecasts, we similarly use the respective IMF point forecast for the target year as an additional predictor in the ARX-model. We additionally augment the model with the IMF point forecast for the current year, such that the model has knowledge of an estimate of the outcome in the current year.444as it does in the AR-annual multi-step predictions, where the model’s estimate for the current year is ”fed back” into the model when making predictions for the next year.
@@
$y_{t+1}=\nu+\alpha\leavevmode\nobreak\ y_{t-1}+\gamma_{1}\leavevmode\nobreak\ \hat{y}_{\text{IMF},{t+1},\text{FN}}+\gamma_{2}\leavevmode\nobreak\ \hat{y}_{\text{IMF},t,\text{FC}}+u_{t+1}.$
@@
$y_{t+1}=\nu+\alpha\leavevmode\nobreak\ y_{t-1}+\gamma_{1}\leavevmode\nobreak\ \hat{y}_{\text{IMF},{t+1},\text{SN}}+\gamma_{2}\leavevmode\nobreak\ \hat{y}_{\text{IMF},t,\text{SC}}+u_{t+1}.$
@@ B.6 Additional Forecast Intervals and Metrics
Figure 12: Empirical quantile coverage deviation in evaluation period 2013-2023, for quantile forecasts based on respective point forecast errors (top panel) and direct quantile forecasts (bottom panel), where forecast quantiles are taken from the respective model’s parametric forecast distribution. Note that “IMF” is generally based on point forecast errors, and is shown in both panels to enable comparison. Empirical quantile coverage is the proportion of realized values that fall below the forecast quantile associated with a given nominal level. Coverage deviation is then computed as the empirical coverage level minus the nominal coverage level. Coverage levels are aggregated over the country and horizon dimension, and reported separately for the two targets (left: GDP growth, right: inflation). The black horizontal line indicates nominal, that is, optimal coverage deviation levels.
@@
1){1)}1 ) Sample-based Continuous Ranked Probability Score, see Jordan et al. (2019).
2){2)}2 ) Different Versions of the Interval Score. $\text{IS}_{W}$ represents a weighted sum of interval scores at extended set of levels $\tau\in\{0.1,0.2,...0.9\}$, with weights as given by Bracher et al. (2021a). $\text{IS}_{W,b}$ represents the analogous weighted sum of interval scores with forecast intervals $\tau\in\{0.5,0.8\}$ from main analysis. $\text{IS}_{U}$ represents a simple equally-weighted average of interval scores at extended set of levels $\tau\in\{0.1,0.2,...0.9\}$. $\text{IS}_{\text{CISS}}$ represents the weighted sum of interval scores with forecast intervals $\tau\in\{0.5,0.8\}$, for a subset of the data that excludes Canada and Japan due to availability of the CISS-data.
3){3)}3 ) Different autoregressive model specifications. ‘AR’ refers to the AR model from main analysis. For more details, please refer to Section B.5 in the supplement.
4){4)}4 ) Different Bayesian vector autoregressive model specifications. ‘BVAR’ refers to the BVAR model from main analysis. For more details, please refer to Section B.5 in the supplement.
5){5)}5 ) Forecast quantiles are taken directly from the parametric forecast distribution of the respective model.
6){6)}6 ) Quantile-wise simple average of models from main analysis, that is, the ’IMF’, ’AR’, and ’BVAR’ forecasts.
@@
1){1)}1 ) Sample-based Continuous Ranked Probability Score, see Jordan et al. (2019).
2){2)}2 ) Different Versions of the Interval Score. $\text{IS}_{W}$ represents a weighted sum of interval scores at extended set of levels $\tau\in\{0.1,0.2,...0.9\}$, with weights as given by Bracher et al. (2021a). $\text{IS}_{W,b}$ represents the analogous weighted sum of interval scores with forecast intervals $\tau\in\{0.5,0.8\}$ from main analysis. $\text{IS}_{U}$ represents a simple equally-weighted average of interval scores at extended set of levels $\tau\in\{0.1,0.2,...0.9\}$. $\text{IS}_{\text{CISS}}$ represents the weighted sum of interval scores with forecast intervals $\tau\in\{0.5,0.8\}$, for a subset of the data that excludes Canada and Japan due to availability of the CISS-data.
3){3)}3 ) Different autoregressive model specifications. ‘AR’ refers to the AR model from main analysis. For more details, please refer to Section B.5 in the supplement.
4){4)}4 ) Different Bayesian vector autoregressive model specifications. ‘BVAR’ refers to the BVAR model from main analysis. For more details, please refer to Section B.5 in the supplement.
5){5)}5 ) Forecast quantiles are taken directly from the parametric forecast distribution of the respective model.
6){6)}6 ) Quantile-wise simple average of models from main analysis, that is, the ’IMF’, ’AR’, and ’BVAR’ forecasts.
@@
1){1)}1 ) Different Versions of the Interval Score. $\text{IS}_{30}$, $\text{IS}_{50}$, $\text{IS}_{80}$ and $\text{IS}_{90}$ represent the interval score for the $30\%$, $50\%$, $80\%$ and $90\%$ intervals, respectively. Note that for the non-direct forecast intervals, IS values shown here are sometimes slightly different than those shown in Table 4 showing results from the main analysis, due to the PAVA-correction algorithm correcting more instances than in the main analysis.
2){2)}2 ) Forecast quantiles are taken directly from the parametric forecast distribution of the respective model.
@@ B.7.1 Inflation
1){1)}1 ) Sample-based Continuous Ranked Probability Score, see Jordan et al. (2019).
2){2)}2 ) Different Versions of the Interval Score. $\text{IS}_{50}$ and $\text{IS}_{80}$ represent the Interval Score for the $50\%$ and $80\%$ intervals, respectively. $\text{IS}_{W,b}$ represents a weighted sum of interval scores from quantile forecasts at base levels from main analysis $\tau\in\{0.5,0.8\}$, with weights as given by Bracher et al. (2021a). $\text{IS}_{U}$ represents a simple equally-weighted average of interval scores, here at base levels $\tau\in\{0.5,0.8\}$.
3){3)}3 ) Forecast quantiles are taken directly from the parametric forecast distribution of the respective model.
@@ B.7.2 GDP Growth
1){1)}1 ) Sample-based Continuous Ranked Probability Score, see Jordan et al. (2019).
2){2)}2 ) Different Versions of the Interval Score. $\text{IS}_{50}$ and $\text{IS}_{80}$ represent the Interval Score for the $50\%$ and $80\%$ intervals, respectively. $\text{IS}_{W,b}$ represents a weighted sum of interval scores from quantile forecasts at base levels from main analysis $\tau\in\{0.5,0.8\}$, with weights as given by Bracher et al. (2021a). $\text{IS}_{U}$ represents a simple equally-weighted average of interval scores, here at base levels $\tau\in\{0.5,0.8\}$.
3){3)}3 ) Forecast quantiles are taken directly from the parametric forecast distribution of the respective model.
@@ B.8.1 GDP Growth
1){1)}1 ) Different Versions of the Interval Score. $\text{IS}_{50}$ and $\text{IS}_{80}$ represent the interval score for the $50\%$ and $80\%$ intervals, respectively. $\text{IS}_{W,b}$ represents the weighted sum of interval scores from quantile forecasts at base levels from main analysis $\tau\in\{0.5,0.8\}$. $\text{IS}_{U}$ represents a simple equally-weighted average of interval scores at the same levels.
2){2)}2 ) Coverage deviation at nominal interval coverage levels $50\%$ and $80\%$. Negative values represent cases where the respective intervals covered fewer observations than the nominal level would indicate, and vice versa for positive values.
@@ B.8.2 Inflation
1){1)}1 ) Different Versions of the Interval Score. $\text{IS}_{50}$ and $\text{IS}_{80}$ represent the interval score for the $50\%$ and $80\%$ intervals, respectively. $\text{IS}_{W,b}$ represents the weighted sum of interval scores from quantile forecasts at base levels from main analysis $\tau\in\{0.5,0.8\}$. $\text{IS}_{U}$ represents a simple equally-weighted average of interval scores at the same levels.
2){2)}2 ) Coverage deviation at nominal interval coverage levels $50\%$ and $80\%$. Negative values represent cases where the respective intervals covered fewer observations than the nominal level would indicate, and vice versa for positive values.

### v4

@@ B.5.2 AR-annual and ARX-annual
$y_{t}=\nu+\alpha\penalty 10000\ y_{t-1}+u_{t},$
where $y_{t}$ is the annual observation (GDP growth rate or inflation rate) of interest, $\nu$ is an intercept term, $1<\alpha<1$ is an autoregressive term and $u_{t}$ is the model’s error term that is assumed to be independently normally distributed with variance $\sigma^{2}$. We estimate $\nu$ and $\alpha$ via ordinary least squares and $\sigma^{2}$ by the mean of squared residuals.
To describe ARX-annual, let $\hat{y}_{\text{IMF},t,\text{FC}}$ denote the IMF point forecast for target year $t$ and target variable $y$, issued in fall of the same year (forecast horizon ”Fall, Current”). Similarly, $\hat{y}_{\text{IMF},t,\text{SC}}$, $\hat{y}_{\text{IMF},t,\text{FN}}$ and $\hat{y}_{\text{IMF},t,\text{SN}}$ respectively denote the point forecast for target year $t$, made in spring of the same year (forecast horizon ”Spring, Current”), fall of the previous year (forecast horizon ”Fall, Next”) and spring of the previous year (forecast horizon ”Spring, Next”). Note thus again that $t$ refers to the target year: for instance, $\hat{y}_{\text{IMF},t,\text{FC}}$ is issued during the same year (year $t$), while $\hat{y}_{\text{IMF},t,\text{FN}}$ is issued in year t1t-1. Both issue forecasts for the same target year.
@@
$y_{t}=\nu+\alpha y_{t-1}+\gamma\penalty 10000\ \hat{y}_{\text{IMF},t,\text{FC}}+u_{t},$
@@
$y_{t}=\nu+\alpha y_{t-1}+\gamma\penalty 10000\ \hat{y}_{\text{IMF},t,\text{SC}}+u_{t}.$
For next-year forecasts, we similarly use the respective IMF point forecast for the target year as an additional predictor in the ARX-model. We additionally augment the model with the IMF point forecast for the current year, such that the model has knowledge of an estimate of the outcome in the current year.555as it does in the AR-annual multi-step predictions, where the model’s estimate for the current year is ”fed back” into the model when making predictions for the next year.
@@
$y_{t+1}=\nu+\alpha\penalty 10000\ y_{t-1}+\gamma_{1}\penalty 10000\ \hat{y}_{\text{IMF},{t+1},\text{FN}}+\gamma_{2}\penalty 10000\ \hat{y}_{\text{IMF},t,\text{FC}}+u_{t+1}.$
@@
$y_{t+1}=\nu+\alpha\penalty 10000\ y_{t-1}+\gamma_{1}\penalty 10000\ \hat{y}_{\text{IMF},{t+1},\text{SN}}+\gamma_{2}\penalty 10000\ \hat{y}_{\text{IMF},t,\text{SC}}+u_{t+1}.$
@@ B.6 Additional Forecast Intervals and Metrics
Figure 12: Empirical quantile coverage deviation in evaluation period 2013-2023, for quantile forecasts based on respective point forecast errors (top panel) and direct quantile forecasts (bottom panel), where forecast quantiles are taken from the respective model’s parametric forecast distribution. Note that “IMF” is generally based on point forecast errors, and is shown in both panels to enable comparison. Empirical quantile coverage is defined as the proportion of observations that fall below the forecast quantiles associated with a given nominal level. Coverage deviation is then computed as the empirical coverage level minus the nominal coverage level. Coverage levels are aggregated over the country and horizon dimension, and reported separately for the two targets (left: GDP growth, right: inflation). The horizontal line indicates nominal, that is, optimal coverage deviation levels.
@@
1){1)} Sample-based Continuous Ranked Probability Score, see Jordan et al. (2019).
2){2)} Different Versions of the Interval Score. $\text{IS}_{W}$ represents a weighted sum of interval scores at extended set of levels $\tau\in\{0.1,0.2,...0.9\}$, with weights as given by Bracher et al. (2021a). $\text{IS}_{W,b}$ represents the analogous weighted sum of interval scores with forecast intervals $\tau\in\{0.5,0.8\}$ from main analysis. $\text{IS}_{U}$ represents a simple equally-weighted average of interval scores at extended set of levels $\tau\in\{0.1,0.2,...0.9\}$. $\text{IS}_{\text{CISS}}$ represents the weighted sum of interval scores with forecast intervals $\tau\in\{0.5,0.8\}$, for a subset of the data that excludes Canada and Japan due to availability of the CISS-data.
3){3)} Different autoregressive model specifications. ‘AR’ refers to the AR model from main analysis. For more details, please refer to Section B.5 in the supplement.
4){4)} Different Bayesian vector autoregressive model specifications. ‘BVAR’ refers to the BVAR model from main analysis. For more details, please refer to Section B.5 in the supplement.
5){5)} Forecast quantiles are taken directly from the parametric forecast distribution of the respective model.
6){6)} Quantile-wise simple average of models from main analysis, that is, the ’IMF’, ’AR’, and ’BVAR’ forecasts.
@@
1){1)} Sample-based Continuous Ranked Probability Score, see Jordan et al. (2019).
2){2)} Different Versions of the Interval Score. $\text{IS}_{W}$ represents a weighted sum of interval scores at extended set of levels $\tau\in\{0.1,0.2,...0.9\}$, with weights as given by Bracher et al. (2021a). $\text{IS}_{W,b}$ represents the analogous weighted sum of interval scores with forecast intervals $\tau\in\{0.5,0.8\}$ from main analysis. $\text{IS}_{U}$ represents a simple equally-weighted average of interval scores at extended set of levels $\tau\in\{0.1,0.2,...0.9\}$. $\text{IS}_{\text{CISS}}$ represents the weighted sum of interval scores with forecast intervals $\tau\in\{0.5,0.8\}$, for a subset of the data that excludes Canada and Japan due to availability of the CISS-data.
3){3)} Different autoregressive model specifications. ‘AR’ refers to the AR model from main analysis. For more details, please refer to Section B.5 in the supplement.
4){4)} Different Bayesian vector autoregressive model specifications. ‘BVAR’ refers to the BVAR model from main analysis. For more details, please refer to Section B.5 in the supplement.
5){5)} Forecast quantiles are taken directly from the parametric forecast distribution of the respective model.
6){6)} Quantile-wise simple average of models from main analysis, that is, the ’IMF’, ’AR’, and ’BVAR’ forecasts.
@@
1){1)} Different Versions of the Interval Score. $\text{IS}_{30}$, $\text{IS}_{50}$, $\text{IS}_{80}$ and $\text{IS}_{90}$ represent the interval score for the $30\%$, $50\%$, $80\%$ and $90\%$ intervals, respectively. Note that for the non-direct forecast intervals, IS values shown here are sometimes slightly different than those shown in Table 4 showing results from the main analysis, due to the PAVA-type correction algorithm correcting more instances than in the main analysis.
2){2)} Forecast quantiles are taken directly from the parametric forecast distribution of the respective model.
@@ B.7.1 Inflation
1){1)} Sample-based Continuous Ranked Probability Score, see Jordan et al. (2019).
2){2)} Different Versions of the Interval Score. $\text{IS}_{50}$ and $\text{IS}_{80}$ represent the Interval Score for the $50\%$ and $80\%$ intervals, respectively. $\text{IS}_{W,b}$ represents a weighted sum of interval scores from quantile forecasts at base levels from main analysis $\tau\in\{0.5,0.8\}$, with weights as given by Bracher et al. (2021a). $\text{IS}_{U}$ represents a simple equally-weighted average of interval scores, here at base levels $\tau\in\{0.5,0.8\}$.
3){3)} Forecast quantiles are taken directly from the parametric forecast distribution of the respective model.
@@ B.7.2 GDP Growth
1){1)} Sample-based Continuous Ranked Probability Score, see Jordan et al. (2019).
2){2)} Different Versions of the Interval Score. $\text{IS}_{50}$ and $\text{IS}_{80}$ represent the Interval Score for the $50\%$ and $80\%$ intervals, respectively. $\text{IS}_{W,b}$ represents a weighted sum of interval scores from quantile forecasts at base levels from main analysis $\tau\in\{0.5,0.8\}$, with weights as given by Bracher et al. (2021a). $\text{IS}_{U}$ represents a simple equally-weighted average of interval scores, here at base levels $\tau\in\{0.5,0.8\}$.
3){3)} Forecast quantiles are taken directly from the parametric forecast distribution of the respective model.
@@ B.8.1 GDP Growth
1){1)} Different Versions of the Interval Score. $\text{IS}_{50}$ and $\text{IS}_{80}$ represent the interval score for the $50\%$ and $80\%$ intervals, respectively. $\text{IS}_{W,b}$ represents the weighted sum of interval scores from quantile forecasts at base levels from main analysis $\tau\in\{0.5,0.8\}$. $\text{IS}_{U}$ represents a simple equally-weighted average of interval scores at the same levels.
2){2)} Coverage deviation at nominal interval coverage levels $50\%$ and $80\%$. Negative values represent cases where the respective intervals covered fewer observations than the nominal level would indicate, and vice versa for positive values.
@@ B.8.2 Inflation
1){1)} Different Versions of the Interval Score. $\text{IS}_{50}$ and $\text{IS}_{80}$ represent the interval score for the $50\%$ and $80\%$ intervals, respectively. $\text{IS}_{W,b}$ represents the weighted sum of interval scores from quantile forecasts at base levels from main analysis $\tau\in\{0.5,0.8\}$. $\text{IS}_{U}$ represents a simple equally-weighted average of interval scores at the same levels.
2){2)} Coverage deviation at nominal interval coverage levels $50\%$ and $80\%$. Negative values represent cases where the respective intervals covered fewer observations than the nominal level would indicate, and vice versa for positive values.